# Design and Control of a Small Humanoid Equipped with Flight Unit and Wheels for Multimodal Locomotion

### 06 Mar 2024 v2

@@ I INTRODUCTION
Several humanoids exist with terrestrial locomotion mechanisms other than bipedal walking such as wheels [6] or crawlers [7]. Compared to walking by bipedal legs, these locomotion methods are based on wheel rotational motion, which makes it easier to achieve high-speed locomotion. In [8], wheels were deployed to a drone for terrestrial locomotion by aerial robots. Thrust can be applied to stabilize the motion of the humanoid presented in [9] and enables aerial locomotion. Therefore, we propose a humanoid equipped with wheels and a flight unit to achieve rapid terrestrial and aerial locomotion. Regarding the humanoid presented in [6] and [7], the employed wheels and crawlers required additional actuators to drive them; however, an increase in weight due to these attached mechanisms severely limits aerial locomotion. Hence, we propose the adoption of passive wheels, which enable a minimum increase in weight and utilize thrust for propulsion.
@@ II-A Flight Unit
$\bm{I}\dot{\bm{\omega}}{\color[rgb]{0,0,0}{}_{\{CoG\}}}=\bm{\tau}-\bm{\omega}{\color[rgb]{0,0,0}{}_{\{CoG\}}}\crossproduct\bm{I}\bm{\omega}{\color[rgb]{0,0,0}{}_{\{CoG\}}},$ (2)
where $M$ denotes the mass of the robot. In addition, $\bm{r}$, $\bm{g}$, $\bm{\omega}{\color[rgb]{0,0,0}{}_{\{CoG\}}}$, $\bm{f}$, $\bm{\tau}$ $\in\mathbb{R}^{3}$ represent the position of the CoG frame w.r.t the world frame, gravitational vector, angular velocity of the CoG frame, force and torque generated by the thrusts in the CoG frame, respectively. $\bm{R}$, $\bm{I}$ $\in\mathbb{R}^{3\times 3}$ denote the rotation matrix of the CoG frame w.r.t the world frame and inertia matrix of the body, respectively. $\bm{f}$ and $\bm{\tau}$ are described in the CoG frame. From Fig. 2, the force and torque in the CoG frame by vectoring thrust force $\bm{\lambda}\in\mathbb{R}^{6}$ can be described as follows using $\bm{Q}\in\mathbb{R}^{6\times 6}$
${\color[rgb]{0,0,0}\begin{bmatrix}\bm{f}\\ \bm{\tau}\end{bmatrix}=\bm{Q}\begin{bmatrix}\lambda_{1,\perp}&\lambda_{1,\parallel}&\cdots&\lambda_{3,\perp}&\lambda_{3,\parallel}\end{bmatrix}^{T}=\bm{Q}\bm{\lambda},}$ (3)
@@ II-D Optimal Orientation of Flight Unit
Figure 5: Plot of feasible torque around pitch axis when tilting by pitch angle. Blue line: $\text{max}\quantity(\tau_{y})$ in (8). Orange line: $\text{min}\quantity(\tau_{y})$ in (8). The range of feasible torque is maximized when the pitch angle of the body is 0 $\mathrm{r}\mathrm{a}\mathrm{d}$.
@@ II-E Desired Clutch Angle
${\color[rgb]{0,0,0}\theta_{clutch}\!\in\!\quantity{\theta_{torso}\quantity(\bm{q})-\theta_{flight\_unit}\mid CoG\quantity(\bm{q})\in S\quantity(\bm{q})},}$ (21)
@@ III-A1 Attitude Control
$\bm{\tau}^{des}=$ $\bm{\omega}{\color[rgb]{0,0,0}{}_{\{CoG\}}}\times\bm{I}\bm{\omega}{\color[rgb]{0,0,0}{}_{\{CoG\}}}+\bm{K}_{\eta,p}\bm{\eta}^{err}$
@@ III-B Leg Locomotion
$\psi^{des}$ $\leftarrow\psi_{torso}\quantity(\bm{q}){\color[rgb]{0,0,0},}$ (27)
@@
$\alpha_{stable}<{\color[rgb]{0,0,0}\ddot{r}}_{z}<g.$ (28)
@@ III-C Wheel Locomotion
$\beta_{stable}<{\color[rgb]{0,0,0}\ddot{r}_{z}}<g,$ (31)

### 09 Mar 2024 v3

@@ I INTRODUCTION
Several humanoids exist with terrestrial locomotion mechanisms other than bipedal walking such as wheels [6] or crawlers [7]. Compared to walking by bipedal legs, these locomotion methods are based on wheel rotational motion, which makes it easier to achieve high-speed locomotion. In [8], wheels were deployed to a drone for terrestrial locomotion by aerial robots. Thrust can be applied to stabilize the motion of the humanoid presented in [9] and enables aerial locomotion. Therefore, we propose a humanoid equipped with wheels and a flight unit to achieve rapid terrestrial and aerial locomotion. Regarding the humanoid presented in [6] and [7], the employed wheels and crawlers required additional actuators to drive them  however, an increase in weight due to these attached mechanisms severely limits aerial locomotion. Hence, we propose the adoption of passive wheels, which enable a minimum increase in weight and utilize thrust for propulsion.
@@ II-A Flight Unit
$\bm{I}\dot{\bm{\omega}}_{\{CoG\}}=\bm{\tau}-\bm{\omega}_{\{CoG\}}\crossproduct\bm{I}\bm{\omega}_{\{CoG\}},$ (2)
where $M$ denotes the mass of the robot. In addition, $\bm{r}$, $\bm{g}$, $\bm{\omega}_{\{CoG\}}$, $\bm{f}$, $\bm{\tau}$ $\in\mathbb{R}^{3}$ represent the position of the CoG frame w.r.t the world frame, gravitational vector, angular velocity of the CoG frame, force and torque generated by the thrusts in the CoG frame, respectively. $\bm{R}$, $\bm{I}$ $\in\mathbb{R}^{3\times 3}$ denote the rotation matrix of the CoG frame w.r.t the world frame and inertia matrix of the body, respectively. $\bm{f}$ and $\bm{\tau}$ are described in the CoG frame. From Fig. 2, the force and torque in the CoG frame by vectoring thrust force $\bm{\lambda}\in\mathbb{R}^{6}$ can be described as follows using $\bm{Q}\in\mathbb{R}^{6\times 6}$
$\begin{bmatrix}\bm{f}\\ \bm{\tau}\end{bmatrix}=\bm{Q}\begin{bmatrix}\lambda_{1,\perp}&\lambda_{1,\parallel}&\cdots&\lambda_{3,\perp}&\lambda_{3,\parallel}\end{bmatrix}^{T}=\bm{Q}\bm{\lambda},$ (3)
@@ II-D Optimal Orientation of Flight Unit
Figure 5: Plot of feasible torque around pitch axis when tilting by pitch angle. Blue line: $\text{max}\quantity(\tau_{y})$ in (8). Orange line: $\text{min}\quantity(\tau_{y})$ in (8). The range of feasible torque is maximized when the pitch angle of the body is 0 $\mathrm{r}\mathrm{a}\mathrm{d}$
Figure 5: Plot of feasible torque around pitch axis when tilting by pitch angle. Blue line: $\text{max}\quantity(\tau_{y})$ in (8). Orange line: $\text{min}\quantity(\tau_{y})$ in (8). The range of feasible torque is maximized when the pitch angle of the body is 0 $\mathrm{r}\mathrm{a}\mathrm{d}$
@@ II-E Desired Clutch Angle
$\theta_{clutch}\!\in\!\quantity{\theta_{torso}\quantity(\bm{q})-\theta_{flight\_unit}\mid CoG\quantity(\bm{q})\in S\quantity(\bm{q})},$ (21)
@@ III-A1 Attitude Control
$\bm{\tau}^{des}=$ $\bm{\omega}_{\{CoG\}}\times\bm{I}\bm{\omega}_{\{CoG\}}+\bm{K}_{\eta,p}\bm{\eta}^{err}$
@@ III-B Leg Locomotion
$\psi^{des}$ $\leftarrow\psi_{torso}\quantity(\bm{q}),$ (27)
@@
$\alpha_{stable}<\ddot{r}_{z}<g.$ (28)
@@ III-C Wheel Locomotion
$\beta_{stable}<\ddot{r}_{z}<g,$ (31)